OpenStudy (anonymous):

Evaluate the expression below when x = 9. 3x + 2x2 - 9 Select one of the options below as your answer: A. 54 B. 180 C. 342 D. 4,365

OpenStudy (bahrom7893):

3*9+2*9^2 - 9 = 27 + 2*81 - 9 = 18 + 2*81 = 18 + 162 = 180 B